OpenStudy (mm3591):

Solve 5p + 9c = p for c

OpenStudy (anonymous):

@mm3591 use this website this helps me a lot (: https://mathway.com/

OpenStudy (mm3591):

Thanks! I'll take a look at it!

OpenStudy (mm3591):

It helped me so much @Moo_Moo17

OpenStudy (usukidoll):

so you understand that we need c to be by itself since it's asking solve 5p+9c=p for c right?

OpenStudy (usukidoll):

While sites like wolfram alpha and mathway are useful, you should also understand the concepts and how to solve these problems manually. I highly doubt you will get an open book open notes exam To solve this manually, subtract 5p from both sides and divide by 9. 5p-5p+9c=p-5p 9c=-4p c=-4p/9

OpenStudy (usukidoll):

plug c back into the equation to see if your left side of the equation = right side of the equation... that's one way to know that you have the correct answer